Write a Systems of Equations for the sentence below:
"A rectangular parking lot has a perimeter of 310 meters. The length is 5 meters more than 4 times the width. What is the length and width?"
Thank you in advance.

Answer by checkley77(12844)

L=4W+5
310=2L+2W
310=2(4W+5)+2W
310=8W+10+2W
310-10=10W
10W=300
W=300/10
W=30 ANS.
L=4*30+5
L=120+5
L=125
PROOF:
310=2*125+2*30
310=250+60
310=310